# Replica Thermodynamic Trade-off Relations: Entropic Bounds on Network Diffusion and Trajectory Observables

### v1

@@ Results
Figure 2: Illustration of (a) the network diffusion bound [Eq. (16)] and (b) the trajectory observable bound [Eq. (17)]. (a) Diffusion in a random walk on a network. A random walker starts from a node in the network, which is represented by dark red. At time $t=\tau$, the random walker diffuses over the network and the probability $p_{\mu}(\tau)$ of the walker position is expressed by red, where darker color corresponds to higher probability. The network diffusion bound provides upper bound to the Rényi entropy of $p_{\mu}(\tau)$. (b) Trajectory observable bound in a Markov process. For the Rényi entropy bound for trajectory observables, we consider a trajectory $\Gamma)$ and its associated observable $N(\Gamma)$. The Rényi entropy is calculated for the distribution $N(\Gamma)$.
@@
As mentioned above, by considering specific observables for $N^{(\otimes K)}$ in Eq. (14), we derive trade-off relations for a single process. In particular, this paper focuses on the Rényi entropy. For an arbitrary probability distribution $p_{n}$, the Rényi entropy with order $\alpha$ ($0<\alpha<\infty$) is given by
$H_{\alpha}[p_{n}]\equiv\frac{1}{1-\alpha}\ln\left(\sum_{n}p_{n}^{\alpha}\right).$ (15)
@@
The Rényi entropy is a one-parameter generalization of several entropies. The real parameter $\alpha$ is called the order of the Rényi entropy. By varying $\alpha$, we can explore different aspects of uncertainty of the distribution. The Rényi entropies form a one-parameter family that continuously interpolate between several entropic definitions. In the limit $\alpha\to 0$, $H_{\alpha}$ approaches the max-entropy, which depends only on the number of outcomes with nonzero probability. In the limit $\alpha\to 1$, $H_{\alpha}$ reduces to the usual Shannon entropy. For $\alpha=2$, the Rényi entropy reduces to the collision entropy. In the limit $\alpha\to\infty$, $H_{\alpha}$ converges to the min-entropy, which is determined solely by the highest probability.
Entropic bound on network diffusion
As the first application of Eq. (14), we consider a random walker on a network, whose dynamics is governed by a continuous time Markov process [Eq. (3)], which is extensively investigated in the literature [46, 47, 48]. It has many applications in web search engines [49], community detection [50], embedding algorithms [51], and computational biology [52], to name but a few. We consider a continuous time random walk starting from the state $B_{\mu_{0}}$. Here, we aim to quantify the extent of diffusion on the network (Fig. 2(a)). Typically, the mean squared displacement, mixing time, or spectral gap of the generator are used [46, 48]. Entropy is another metric used to quantify diffusion in a network. The diversity entropy [53, 54] is defined as the entropy of the walker’s position probabilities and measures how widely a random walker can spread over the network. The entropy rate [55] quantifies how quickly uncertainty increases as the random walk evolves.
Let $p_{\mu}(\tau\mid\mu_{0})$ be the probability of being in $B_{\mu}$ at time $t=\tau$ starting from the state $B_{\mu_{0}}$ at time $t=0$. In this case, the extent of diffusion from node $\mu_{0}$ can be quantified by the Rényi entropy $H_{\alpha}[p_{\mu}(\tau\mid\mu_{0})]$. When this Rényi entropy is zero, it indicates that no propagation has occurred in the network, and as the degree of propagation increases, the value of this Rényi entropy increases. Using Eq. (14), the following relation holds for the Rényi entropy (see Methods):
$H_{\alpha}[p_{\mu}(\tau|\mu_{0})]\leq\frac{\alpha\tau}{\alpha-1}\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}\,\,\,(\alpha=2,3,\cdots),$ (16)
which is the second result of this study. The left-hand side of Eq. (16) represents the Rényi entropy and it is bounded from above by the $\tau$ and the local escape rate from the initial node $B_{\mu_{0}}$, which is the sum of the transition rates from the initial node $B_{\mu_{0}}$ to the nodes that can be reached within one jump. Note that $\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}=\mathfrak{a}(t=0)$ corresponds to the dynamical activity at the initial time. Surprisingly, this right-hand side depends only on the time $\tau$ and on local escape rate from $B_{\mu_{0}}$ (see Fig. 3). In other words, regardless of how large the overall network is, we can compute the upper bound of the Rényi entropy using only information about the initial node and the transition rates to nodes that are reachable in a single hop. For integer values of $\alpha$, $\alpha\in\{2,3,\cdots\}$, the relation is exact. Since the inequality of Eq. (16) still holds for real value of $\alpha>1$, we conjecture that this relation holds for arbitrary $\alpha>1$.
Entropic bound on trajectory observable
Next, we consider the uncertainty in trajectory observables. Trajectory observables play a fundamental role in uncertainty relations in stochastic thermodynamics (Fig. 2(b)). Consider a trajectory observable $N(\Gamma)$, which vanishes for $\Gamma_{\varnothing}$ [Eq. (11)]; that is, when there is no jump within $0,\tau]$, the observable $N(\Gamma)$ vanishes. This is a generalization of counting observables often considered in stochastic thermodynamics. From Eq. (14), we obtain the following relation (see Methods):
$H_{\alpha}[P(N)]\leq\frac{\alpha\tau}{\alpha-1}\mathfrak{a}(t=0)\,\,\,\,(\alpha=2,3,\cdots),$ (17)
@@
which is the third result of this study. The left-hand side of Eq. (17) is the Rényi entropy of the observable $N$, while the right-hand side depends on the dynamical activity and on the duration $\tau$. For example, consider dynamics with no jumps at all, or the case $\tau=0$. Since there are no jumps, the distribution of $N$ is $P(N=0)=1$, and the entropy of $N(\Gamma)$ is zero. In this case, the right-hand side also vanishes, so the result is consistent. On the other hand, for dynamics with a very large number of jumps, the distribution acquires weight at many different values of $N$, and the entropy increases. In fact, the right-hand side increases with both the dynamical activity and the duration $\tau$, which matches our intuition. Note also that for the counting observable $N$, there is no upper bound on its possible values, and therefore there is no upper bound on the entropy either. For integer values of $\alpha$, $\alpha\in\{2,3,\cdots\}$, the relation is exact. Since the inequality of Eq. (17) formally holds for the real value of $\alpha>1$, again, we conjecture that this relation holds for arbitrary $\alpha>1$.
@@
Traditionally, uncertainty in trajectory observables has been characterized using the variance $\mathrm{Var}[N]$, whose relative value is known to be bounded from below by the inverse of the dynamical activity. For a steady-state Markov process and the counting observable $N_{\mathrm{c}}$, it is known that the following relation holds:
$\frac{\mathrm{Var}[N_{\mathrm{c}}]}{\mathbb{E}[N_{\mathrm{c}}]^{2}}\geq\frac{1}{\mathfrak{a}_{\mathrm{ss}}\tau},$ (18)
@@
where $\mathfrak{a}_{\mathrm{ss}}$ is the stead-state dynamical activity. Our results show that this fundamental constraint extends beyond relative variance: the uncertainty measure based on Rényi entropy is also limited by the dynamical activity. However, there is an important difference between Eq. (17) and Eq. (18). When the variance $\mathrm{Var}[N]$ decreases, the uncertainty in $N$ becomes smaller; thus Eq. (18) provides a lower bound on the uncertainty in $N$. By contrast, Eq. (17) yields an upper bound on the uncertainty in $N$. Therefore, these two relations give complementary lower and upper bounds on the uncertainty of $N$, each expressed in terms of different quantities. The Rényi entropy reduces to the Shannon entropy for $\alpha\to 1$; however, the bound in Eq. (17) is not defined for $\alpha=1$. Regarding the Shannon entropy for trajectory observables, Ref. [56] derived a lower bound for the Shannon entropy using the entropy production. The bound in Eq. (17) provides the upper bound of the Rényi entropy and the bound comprises the dynamical activity.
So far, we are concerned with the Rényi entropy for probability distributions. We can consider other observables to obtain trade-off relations by using replicas. We consider the variance:
$V^{(\otimes 2)}(\Gamma^{(\otimes 2)})=N^{(1)}(\Gamma^{(1)})^{2}-N^{(1)}(\Gamma^{(1)})N^{(2)}(\Gamma^{(2)}).$ (19)
Using the operator in Eq. (19), its expectation is
$\mathbb{E}\left[V^{(\otimes 2)}(\Gamma^{(\otimes 2)})\right]=\mathbb{E}[N(\Gamma)^{2}]-\mathbb{E}[N(\Gamma)]^{2}=\mathrm{Var}[N(\Gamma)].$ (20)
Assume that the minimum of $N(\Gamma)$ is 0. Substituting Eq. (20) into (14), we obtain the upper bound of the variance as follows:
$\mathrm{Var}[N]\leq N_{\max}^{2}\left(1-e^{-2\tau\mathfrak{a}(t=0)}\right),$ (21)
where $N_{\max}$ is the maximum $N_{\max}\equiv\max_{\Gamma}\left|N(\Gamma)\right|$. Equation (21) gives an upper bound on the observable $N$. To make the discussion concrete, let us focus on the counting observable $N_{\mathrm{c}}$. If no jump occurs, the right-hand side of Eq. (21) vanishes, and the variance of $N_{\mathrm{c}}$ is also zero, so the bound is satisfied. However, for typical counting observables there is no upper limit on the number of jumps, so $N_{\max}=\infty$, and Eq. (21) is no longer applicable. In contrast, when we consider qudits and related finite-dimensional systems, which are often used as working media in heat engines, the observable is naturally bounded. In such cases $N_{\max}$ is finite, and Eq. (21) provides a meaningful bound.
Figure 3: Illustration of the dependence of the Rényi entropic bound. The Rényi entropic bound for the network diffusion depends on $\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}$, where $B_{\mu_{0}}$ is the initial state, the local escape rate from $B_{\mu_{0}}$. The local escape rate from $B_{\mu_{0}}$ is the sum of transition rates to which the nodes reachable from the initial state with one jump. Suppose the initial state is $B_{1}$, which is colored with dark red. Then, the reachable states are represented by purple nodes. Therefore, the local escape rate becomes $\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}=W_{21}+W_{31}+W_{41}$.
Numerical simulation
We perform numerical calculations for the diffusion bound of a random walker in a network [Eq. (16)]. We generate a random network and consider a continuous time random walk on the network. The transition rate $W_{\mu^{\prime}\mu}$ and the initial node are chosen randomly. We calculate the probability distribution at time $t=\tau$, where $\tau$ is determined randomly. The range for these random parameters is shown in the caption of Fig. 4. We repeat this procedure to obtain random realizations and calculate the Rényi entropy (the left-hand side of Eq. (16)) and $\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}$ (the right-hand side of Eq. (16)). For each random realization, we plot $H_{\alpha}[p_{\mu}(\tau)]$ versus $\frac{\alpha\tau}{\alpha-1}\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}$ as a point in Fig. 4, where the dashed line indicates the equality case in Eq. (16). In Figs. 4(a)-(c), we plot the results for (a) $\alpha=1.5$, (b) $\alpha=2$, and (c) $\alpha=3$. For all values of $\alpha$, the points lie above the dashed line, providing numerical verification of Eq. (16). The bound is quite tight, especially for larger $\alpha$. Note that for $\alpha=2$ and $3$, Eq. (16) is exact. However, for a non-integer $\alpha$, the relation has not been proven, so Eq. (16) in $\alpha=1.5$ should be regarded as a conjecture. Our numerical simulations indicate that this conjectured bound also holds in the non-integer case.
Next, we perform numerical calculations for the bound for the trajectory observable [Eq. (17)]. We generate a continuous time Markov process, where the transition rate $W_{\mu^{\prime}\mu}$ and the initial probability distribution $p_{\mu}(0)$ are determined randomly. We consider the counting observable $N_{\mathrm{c}}(\Gamma)$, which simply counts the number of total jumps within the interval $0,\tau]$, where $\tau$ is randomly determined. The range for these random parameters is shown in the caption of Fig. 4. We repeat this procedure to obtain random realizations and calculate the Rényi entropy and the dynamical activity. For each random realization, we plot $H_{\alpha}[P(N)]$ versus $\frac{\alpha\tau\mathfrak{a}(t=0)}{\alpha-1}$ by a point in Fig. 5. In Figs. 5(a)-(c), we plot the results for (a) $\alpha=1.5$, (b) $\alpha=2$, and (c) $\alpha=3$, where the dashed line denotes the equality case of Eq. (17). Again, in all cases of $\alpha$, the points are above the dashed line, which numerically verifies Eq. (17). The derived relation numerically holds for $\alpha=1.5$, which is non-integer and is a conjectured bound.
Quantum generalization

### v2

@@ II.3 Entropic measures
Figure 2: Illustration of two scenarios considered in trade-off relations, (a) the trajectory observable and (b) the state distribution. (a) Trajectory observable bound in a Markov process. For the Tsallis entropy bound for trajectory observables, we consider a trajectory $\Gamma$ and its associated observables $N_{\bullet}(\Gamma)$ and $N_{\circ}(\Gamma)$. The Tsallis entropy is considered for $N_{\bullet}(\Gamma)$ and $N_{\circ}(\Gamma)$. (b) State distribution bound in the diffusion on a network. A random walker starts from a node in the network, which is represented by dark red. At time $t=\tau$, the random walker diffuses over the network and the probability $p_{\mu}(\tau)$ of the walker position is expressed by red, where darker color corresponds to higher probability. The Tsallis entropy is considered for $p_{\mu}(\tau)$.
@@
By considering specific observables for $N_{\bullet}^{(\otimes K)}$ and $N_{\circ}^{(\otimes K)}$ in Eqs. (14) and (17), we derive trade-off relations for a single process. In particular, we focus on one-parameter generalizations of entropy that quantify different aspects of uncertainty in an arbitrary probability distribution $\{p_{n}\}$. The Rényi entropy [46] of order $\alpha$ $(0<\alpha<\infty)$ is defined as
$H_{\alpha}^{\mathrm{R}}[p_{n}]\equiv\frac{1}{1-\alpha}\ln\left(\sum_{n}p_{n}^{\alpha}\right).$ (19)
@@
Varying $\alpha$ continuously interpolates between several standard entropic quantities: in the limit $\alpha\to 0$, $H_{\alpha}^{\mathrm{R}}$ approaches the max-entropy, which depends only on the number of outcomes with nonzero probability; as $\alpha\to 1$, it reduces to the Shannon entropy; for $\alpha=2$, it becomes the collision entropy; and in the limit $\alpha\to\infty$, it converges to the min-entropy determined solely by the largest probability. Another widely used one-parameter generalization is the Tsallis entropy [47, 48], defined by
$H_{q}^{\mathrm{T}}[p_{n}]\equiv\frac{1-\sum_{n}p_{n}^{q}}{q-1},$ (20)
where $q$ is the entropic index. As with Rényi entropy, changing $q$ explores different uncertainty features and connects multiple entropic definitions: in the limit $q\to 0$, $H_{q}^{\mathrm{T}}$ approaches $\mathfrak{W}-1$, where $\mathfrak{W}$ is the number of outcomes with nonzero probability; as $q\to 1$, it reduces to the Shannon entropy; for $q=2$, it yields $H_{2}^{\mathrm{T}}=1-\sum_{n}p_{n}^{2}$, i.e., the complement of the collision probability; and in the limit $q\to\infty$, it tends to 0 for any non-degenerate distribution. For $q>1$, the Tsallis entropy satisfies
$0\leq H_{q}^{\mathrm{T}}<\frac{1}{q-1}\;\;(q>1).$ (21)
Both entropies are related via
$H_{\alpha}^{\mathrm{R}}$ $=\frac{1}{1-\alpha}\ln\left[1+(1-\alpha)H_{\alpha}^{\mathrm{T}}\right],$ (22)
$H_{q}^{\mathrm{T}}$ $=\frac{\exp\left((1-q)H_{q}^{\mathrm{R}}\right)-1}{1-q}.$ (23)
The Rényi entropy is widely used in statistics and quantum information theory. In contrast, the Tsallis entropy is mainly applied to non-extensive statistical mechanics and the study of complex systems.
II.4 Entropic bound on trajectory observable
We consider the uncertainty in trajectory observables. Trajectory observables play a fundamental role in uncertainty relations in stochastic thermodynamics (Fig. 2(a)). Consider a general observable $N_{\bullet}(\Gamma)$. Using Eq. (14), we obtain the upper bound on the time derivative of the Tsallis entropy (see Appendix A for the derivation):
$\left|\frac{\partial}{\partial\tau}H_{q}^{\mathrm{T}}\left[P\left(N_{\bullet}\right)\right]\right|\leq\frac{\sqrt{q\mathcal{A}(\tau)}}{2\tau(q-1)},$ (24)
where $q=2,3,\cdots$. The bound in Eq. (24) provides the upper bound of the change in the Tsallis entropy. Equation (24) shows that the change in entropy is larger for larger $\mathcal{A}(\tau)$. By performing the integration in Eq. (24), we can derive the upper bound on the Tsallis entropy with respect to the observable $N_{\circ}$ (see Appendix A for the derivation):
$H_{q}^{\mathrm{T}}[P(N_{\circ})]\leq\frac{1}{q-1}\sin\left(\frac{1}{2}\int_{0}^{\tau}\frac{\sqrt{q\mathcal{A}(t)}}{t}dt\right)^{2},$ (25)
where $q=2,3,\cdots$. Note that in Eq. (25), the following condition should be satisfied:
$0\leq\frac{1}{2}\int_{0}^{\tau}\frac{\sqrt{q\mathcal{A}(t)}}{t}dt\leq\frac{\pi}{2}.$ (26)
Outside the range of Eq. (26), the bound of Eq. (25) provides a trivial upper bound 1/(q1)1/(q-1), which directly follows from Eq. (21). On the right-hand side of Eq. (25), the integral $\int_{0}^{\tau}dt\,\sqrt{\mathcal{A}(t)}/t$ often appears in thermodynamic trade-off relations [44]. We have demonstrated the consequence of considering Eq. (14) in the replica scenario. We can derive another trade-off relation from Eq. (17). From Eq. (17), we obtain another trade-off relation (see Appendix C for the derivation):
$H_{\alpha}^{\mathrm{R}}[P(N_{\circ})]\leq\frac{\alpha\tau}{\alpha-1}\mathfrak{a}(t=0),$ (27)
@@ II.4 Entropic bound on trajectory observable
where $\alpha=2,3,\cdots$. The left-hand side of Eq. (27) is the Rényi entropy of the observable $N_{\circ}$, while the right-hand side depends on the dynamical activity and the duration $\tau$. For example, consider dynamics with no jumps at all, or the case $\tau=0$. Since there are no jumps, the distribution of $N_{\circ}$ is $P(N_{\circ}=0)=1$, and the entropy of $N_{\circ}(\Gamma)$ is zero. In this case, the right-hand side also vanishes, so the result is consistent. On the other hand, for dynamics with a very large number of jumps, the distribution acquires weight at many different values of $N_{\circ}$, and the entropy increases. In fact, the right-hand side increases with both the dynamical activity and duration $\tau$, which is consistent with our intuition. Equation (27) is equivalently written as the bound for the Tsallis entropy via Eq. (23):
$H_{q}^{\mathrm{T}}[P(N_{\circ})]\leq\frac{1-e^{-q\tau\mathfrak{a}(t=0)}}{q-1}.$ (28)
Both left-hand sides of Eqs. (25) and (28) are Tsallis entropies, and the right-hand sides involve the dynamical activity. In a general transient setting, it is unclear which of the two is tighter; however, in the steady state, Eq. (28) provides the tighter bound.
@@
Traditionally, uncertainty in trajectory observables has been characterized using variance $\mathrm{Var}[N]$, whose relative value is known to be bounded from below by the inverse of the dynamical activity. For a steady-state Markov process and the counting observable $N(\Gamma)$ [Eq. (5)], it is known that the following relation holds [3, 5]:
$\frac{\mathrm{Var}[N]}{\mathbb{E}[N]^{2}}\geq\frac{1}{\mathfrak{a}_{\mathrm{ss}}\tau},$ (29)
where $\mathfrak{a}_{\mathrm{ss}}$ is the steady-state dynamical activity. Moreover, for non-steady-state conditions and for the observable $N_{\circ}(\Gamma)$ satisfying Eq. (13), the following relation is known to hold [45, 49]:
$\frac{\mathrm{Var}[N_{\circ}]}{\mathbb{E}[N_{\circ}]^{2}}\geq\frac{1}{e^{\mathfrak{a}(t=0)\tau}-1},$ (30)
@@
which follows directly from Eq. (17) with $K=1$. Our results show that this fundamental constraint extends beyond relative variance: the uncertainty measure based on Tsallis (or Rényi) entropy is also limited by the dynamical activity. However, there is an important difference between Eq. (28) [or Eq. (25)] and Eq. (29). When the variance $\mathrm{Var}[N]$ decreases, the uncertainty in $N$ becomes smaller; thus Eq. (29) provides a lower bound on the uncertainty in $N$. By contrast, Eq. (28) [or Eq. (25)] yields an upper bound on the uncertainty in the trajectory observable. Therefore, these two relations give complementary lower and upper bounds on the uncertainty of $N$, each expressed in terms of different quantities. The Tsallis entropy reduces to the Shannon entropy for $q\to 1$; however, the bound in Eq. (28) is not defined for $q=1$. Regarding the Shannon entropy for trajectory observables, Ref. [50] derived a lower bound for the Shannon entropy using the entropy production. The bound in Eq. (28) provides the upper bound of the Tsallis entropy and the bound comprises the dynamical activity.
II.5 Entropic bound on state distribution
So far we are concerned with entropic trade-off relations for the trajectory observables $N_{\bullet}(\Gamma)$ and $N_{\circ}(\Gamma)$. Next, we derive entropic trade-off relations for the state probability distribution $p_{\mu}(t)$. The entropy of the state distribution can quantify the extent of diffusion in a network. Specifically, let us consider a random walker on a network whose dynamics is governed by a continuous time Markov process [Eq. (3)], which is extensively investigated in the literature [51, 52, 53]. It has many applications in web search engines [54], community detection [55], embedding algorithms [56], and computational biology [57], to name but a few. Here, our objective is to quantify the extent of diffusion in the network (Fig. 2(b)). Typically, the mean squared displacement, mixing time, or the spectral gap of the generator are used [51, 53]. Entropy is another metric that is used to quantify diffusion in a network. The diversity entropy [58, 59] is defined as the entropy of the probability of position of the walker and measures how widely a random walker can spread over the network. The entropy rate [60] quantifies how quickly the uncertainty increases as the random walk evolves.
The degree of diffusion can be quantified by the Tsallis entropy (or Rényi) $H_{q}^{\mathrm{T}}[p_{\mu}(\tau)]$ (or $H_{\alpha}^{\mathrm{R}}[p_{\mu}(\tau)]$). When this entropy is zero, it indicates that no propagation has occurred in the network, and as the degree of propagation increases, the value of this entropy increases. From Eq. (24), we obtain
$\left|\frac{\partial}{\partial\tau}H_{q}^{\mathrm{T}}\left[p_{\mu}\left(\tau\right)\right]\right|\leq\frac{\sqrt{q\mathcal{A}(\tau)}}{2\tau(q-1)},$ (31)
where $q=2,3,\cdots$. The derivation of Eq. (31) is shown in Appendix A. As expected, a larger dynamical activity allows for faster diffusion, which is quantified by the change in the Tsallis entropy. By integrating Eq. (31), we obtain the upper bound on the Tsallis entropy of the state distribution:
$\left|H_{q}^{\mathrm{T}}[p_{\mu}(\tau)]-H_{q}^{\mathrm{T}}[p_{\mu}(0)]\right|\leq\frac{\sqrt{q}}{2(q-1)}\int_{0}^{\tau}\frac{\sqrt{\mathcal{A}(t)}}{t}dt,$ (32)
where $q=2,3,\cdots$. Equation (32) describes the difference of the Tsallis entropy at $t=0$ and $t=\tau$. The bounds in Eqs. (31) and (32) hold for arbitrary initial state $p_{\mu}(0)$. When the initial state is $B_{\mu_{0}}$ (i.e., $p_{\mu_{0}}(0)=1$), then the tighter relation holds (see Appendix A for the derivation):
$H_{q}^{\mathrm{T}}[p_{\mu}(\tau\mid\mu_{0})]\leq\frac{1}{q-1}\sin\left(\frac{1}{2}\int_{0}^{\tau}\frac{\sqrt{q\mathcal{A}(t)}}{t}dt\right)^{2},$ (33)
where $q=2,3,\cdots$ and the condition given in Eq. (26) should be satisfied. As in the trajectory observable case, we can derive another trade-off relation from Eq. (17). Using Eq. (17), the following relation holds for the Rényi entropy (see Appendix C):
$H_{\alpha}^{\mathrm{R}}[p_{\mu}(\tau\mid\mu_{0})]\leq\frac{\alpha\tau}{\alpha-1}\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}},$ (34)
where $\alpha=2,3,\cdots$. From Eq. (23), Eq. (34) is equivalently written as the bound for the Tsallis entropy:
$H_{q}^{\mathrm{T}}[p_{\mu}(\tau\mid\mu_{0})]\leq\frac{1-e^{-q\tau\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}}}{q-1},$ (35)
where $q=2,3,\cdots$. The left-hand sides of Eqs. (33) and (35) are the same, and in both equations the upper bound of the Tsallis entropy at time $\tau$ is given by the dynamical activity; a larger dynamical activity allows for a larger Tsallis entropy. On the other hand, while Eq. (33) requires information beyond the neighborhood of the initial node, Eq. (35) does not. Specifically, in Eq. (35), it is bounded from above by $\tau$ and the local escape rate from the initial node $B_{\mu_{0}}$, which is the sum of the transition rates from the initial node $B_{\mu_{0}}$ to the nodes that can be reached within a jump ( note that $\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}=\mathfrak{a}(t=0)$ corresponds to the dynamical activity at the initial time). Surprisingly, this right-hand side of Eq. (35) depends only on the time $\tau$ and on the local escape rate from $B_{\mu_{0}}$ (see Fig. 3). In other words, regardless of how large the overall network is, we can compute the upper bound of the entropy using only information about the initial node and the transition rates to nodes that are reachable in a single hop.
II.6 Bound on extreme value
By using replicas, we can consider more general observables. As the next example, we derive a trade-off relation for the extreme-value of a stochastic process. Extreme-value statistics studies the behavior of the largest (and similarly, the smallest) value in a collection of random variables $\{X_{k}\}$; that is, it focuses on the random variable $\max_{k}X_{k}$ (and analogously $\min_{k}X_{k}$):
$X^{M\max}\equiv\max_{1\leq k\leq M}X_{k}.$ (36)
Consider the observable $N_{\circ}^{M,\max}$, which is the maximum of $N_{\circ}^{(k)}(\Gamma^{(k)})$:
$N_{\circ}^{M\max}(\Gamma^{(\otimes M)})\equiv\max_{1\leq k\leq M}N_{\circ}^{(k)}(\Gamma^{(k)}).$ (37)
When the state of the Markov process describes the position, $N^{M\max}_{\circ}$ expresses the maximum position at time $t=\tau$. In reliability engineering, the system load can be modeled using a Markov process [61, 62]. In this case, the maximum value of the stochastic process represents the maximum load of the parallel system. Since $N_{\circ}^{M\max}$ satisfies Eq. (6), using Eq. (18), we obtain
$\frac{\mathrm{Var}[N_{\circ}^{M\max}]}{\mathbb{E}[N_{\circ}^{M\max}]^{2}}\geq\frac{1}{e^{M\mathfrak{a}(t=0)\tau}-1}.$ (38)
Since the dynamical activity is multiplied by $M$, the relative variance of the maximum observable $N_{\circ}^{M\max}$ can be smaller than those of a single process. Usually, when we take the max operation, the mean becomes larger and the variance becomes smaller than those of the original distribution. For example, consider two independent random variables, each following a standard normal distribution. If we take the maximum of these two variables, the resulting variable will have a mean of $1/\sqrt{\pi}$ and a variance of $1-1/\pi$. Its mean is larger, and its variance is smaller when compared to the original standard normal distribution.
Replicas can also be used to establish bounds on the entropy of $N_{\circ}^{M\max}$. We define $K=qM$ replicas, where $q$ is the entropic order and $M$ is the number of observations used to compute the maximum. We then consider the Tsallis entropy of the probability distribution of $N^{M\max}_{\circ}$. Applying the same technique as outlined in Eqs. (24), (25), and (28), we derive the following bounds:
$\left|\frac{\partial}{\partial\tau}H_{q}^{\mathrm{T}}[P(N_{\circ}^{M\max})]\right|\leq\frac{\sqrt{qM\mathcal{A}(\tau)}}{2\tau(q-1)},$ (39)
$H_{q}^{\mathrm{T}}[P(N_{\circ}^{M\max})]\leq\frac{1}{q-1}\sin\left(\frac{1}{2}\int_{0}^{\tau}\frac{\sqrt{qM\mathcal{A}(t)}}{t}dt\right)^{2},$ (40)
and
$H_{q}^{\mathrm{T}}[P(N_{\circ}^{M\max})]\leq\frac{1-e^{-qM\tau\mathfrak{a}(t=0)}}{q-1},$ (41)
where $q=2,3,\cdots$. Again, $0\leq(1/2)\int_{0}^{\tau}\sqrt{qM\mathcal{A}(t)}/t\,dt\leq\pi/2$ should be satisfied in Eq. (40).
II.7 Quantum replica Markov process